OpenStudy (anonymous):

simplify the number using the imaginary unit i square root -28

OpenStudy (anonymous):

i got \[-2\sqrt{-7}\]

OpenStudy (kayliwarden):

\[i ^{2}=-1\] so: (-1)(28) \[i ^{?}\sqrt{28}\] \[i^{2} \times \sqrt{4} \times \sqrt{7}\] \[i ^{2} \times 2\sqrt{7}\] \[-2\sqrt{7}\]

OpenStudy (kayliwarden):

yeah you can't take the square root of a negative number

OpenStudy (anonymous):

my options are \[2\sqrt{-7}\],\[-2\sqrt{-7}\], and \[2i \sqrt{7}\]

OpenStudy (kayliwarden):

hmm then it's probably that last one

OpenStudy (anonymous):

that's correct :3 thank you for your help

OpenStudy (kayliwarden):

anytime